# Investigating MAXI J1752--457 with NuSTAR in the aftermath of a superburst

### v1

@@ 3.2 Timing analysis
We investigated the timing properties of MAXI J1752--457 during the second NuSTAR observation. Beginning with the clean Mode 1 and Mode 06 event files, we first corrected the photon arrival times for the motion of NuSTAR about the barycenter of the Solar System using the ftools utility, barycorr. We specified the source position as measured by Swift/XRT and reported by Liu et al. (2024). We used NuSTAR clockfile 20100101v203, produced on April 15, 2025. Next, we filtered the events using circular source regions with radii $100^{\prime\prime}$ in order to encompass as many source counts as possible given the centroid smearing in Mode 06. Unlike our spectral analysis, and in order to avoid complicated good time intervals (GTIs), we did not split the Mode 06 events into different CHU combinations.
We next used the timing analysis software packages HENDRICS and Stingray (Huppenkothen et al., 2019; Bachetti et al., 2024a, b) to complete the rest of the timing analysis. We found that the background dominates the source spectrum above 30 keV, so we filtered out events with energy below 3 keV and above 30 keV. For each focal plane module, we merged the Mode 1 and Mode 06 event lists. Using Stingray, we converted the event lists into light-curves with bins of $dt=2048^{-1}$ s. We split these light-curves into 16 segments of length $1024$ s each. We calculated the averaged power density spectrum (PDS) over these segments for FPMA and FPMB, then produced a combined PDS which we corrected for deadtime using the Fourier Amplitude Difference normalization described in Bachetti & Huppenkothen (2018). We repeated this procedure in order to produce a background PDS, using only events in background regions with radius $150^{\prime\prime}$, away from the source. The resulting Leahy-normalized PDS, binned logarithmically for clarity, are shown in the left panel of Figure 4, where the source PDS is shown in black, and the background PDS is shown in magenta. In order to investigate the energy-dependence of the source variability, we also calculated the source PDS for events with 3–6 keV and for events with 6–30 keV, shown in green and orange, respectively, in the right panel of Figure 4.
@@
4 Discussion and Conclusions
MAXI and NinjaSat observations of MAXI J1752--457 during a superburst in November, 2024, showed that the flux and spectral evolution were consistent with an accretion-induced superburst (Aoyama et al., 2025). While these observations were sensitive to the blackbody component of the spectrum, and have demonstrated clear evolution of the temperature and radius of the blackbody, NuSTAR’s unique sensitivity above $10$ keV has allowed us to put far better constraints on the blackbody parameters following the rapid decay of the burst. We found that by the time of the second NuSTAR observation, which occured about $79$ hours after the onset of the burst, the blackbody temperature had decreased to $0.60\pm 0.1$ keV, and the inferred radius (assuming a distance of $8$ kpc) had shrunk to $6.0^{+0.4}_{-0.3}$ km.
@@ 4 Discussion and Conclusions
NuSTAR observations also revealed that, in addition to a blackbody component, the source spectrum contained a hard X-ray excess which is well-described by a steep power law with photon index $\Gamma\approx 4$. This indicates the presence of a non-thermal plasma in which soft X-ray photons gain energy via Compton upscattering. Accreting neutron stars at similar luminosities ($10^{35-36}$$\mathrm{erg\,s^{-1}}$) tend to exhibit power law spectral components with photon indices in the range $3.01.5<\Gamma<$ 3.0 (Wijnands et al., 2015). We therefore revisit the spectra obtained during the second NuSTAR observation in order to explore the possibility of a more typical photon index.
@@
As we noted in Section 3.1, when we allowed for a high-energy cutoff to be applied to the power law component, the best-fit value of $E_{\mathrm{cut}}$ was far beyond the NuSTAR energy range, indicating that a power law without a cutoff energy is preferred. However, including a high-energy cutoff allows for more flexibility in the photon index parameter as a low cutoff energy can compensate for a shallower power law. For the cutoff power law model, the $90\%$ confidence interval for the photon index is expanded, giving $\Gamma=3.7^{+0.2}_{-0.7}$. Furthermore, we calculated the $99\%$ confidence intervals for the photon index and cutoff energy by varying each parameter independently until the fit statistic had increased by 6.635. We found intervals of $4.02.3<\Gamma<$ 4.0 and $E_{\mathrm{cut}}>8$ keV. In order to explore the parameter space in even more detail, we stepped over the photon index and the cutoff energy and calculated the $W$-statistic at each step. The resulting grid is shown in Figure 5, with contours representing the 1-$\sigma$, $90\%$, and $99\%$ confidence regions. We again find that the spectra strongly prefer a steep power law with a $90\%$ confidence region that indicates a photon index greater than 2.5 and a cutoff energy above 10 keV. While photon index values of $<2.5$ are allowed within the 99% confidence region, they would require cutoff energies of less than 10 keV. This combination of parameters would imply a very high optical depth (Kara & García, 2025). While we cannot completely rule out a more typical value for the photon index, spectral modeling of MAXI J1752--457 strongly prefers an abnormally steep power law in order account for the hard X-ray excess above about 7 keV.
@@
Existing analyses of the hard X-ray spectra of accreting neutron stars during and after superburst are limited due to the relative infrequency of these bursts. In fact, at the time of writing, we are unable to identify any published observations of superbursts with NuSTAR. However, the Rossi X-ray Timing Explorer (RXTE), thanks to its All-Sky Monitor (ASM) instrument, observed 12 superbursts in the hard X-ray during its lifetime (in’t Zand, 2017). Often, analyses of the spectral evolution of superbursts with RXTE fixed the power law component to the value measured for persistent emission, typically around $\Gamma\approx 2$ (e.g. Keek et al., 2008). In some instances, significant softening of the superburst decay spectrum compared to the persistent spectrum has been observed (e.g. Kuulkers et al., 2004).
@@
Studies of superbursts in the soft X-ray band ($<10$ keV), although insensitive to the hard X-ray emission, have hinted at high photon indices during or shortly after superbursts or similar bursts. For example, Bult et al. (2021) reported on NICER measurements of the spectral evolution of IGR J17062--6143 during an intermediate duration Type I X-ray burst observed in 2020. While not categorized as a superburst, the decay of this burst lasted several days, making it comparable to the case of MAXI J1752--457. These authors reported that, when a relativistic reflection component was included in the spectral model of the decay emission, the photon index reached values in the range of $53<\Gamma<$ 5 throughout the two-day period following the burst before settling to more typical values around $\Gamma\approx 2.5$. Notably, however, when the spectrum was modeled without relativistic reflection, the opposite trend in the power law photon index was recovered, with $\Gamma$ increasing from around 1 to 2 throughout the burst decay.
@@
In a recently published NICER study of the 2021 superburst of 4U 1820--30, Peng et al. (2025) modeled the persistent, burst, and decay emission using a combination of a spherical blackbody and a Comptonization component represented by the Xspec model compTT. The authors noted that during the burst, the contribution of the Comptonization component could be neglected, but they reported that during the first few hours of the decay, they measured a significantly lower electron temperature than during the persistent emission, indicating a softened hard X-ray component. Both of these NICER studies, however, highlight the difficulty of disentangling spectral components with soft X-ray observations alone when the emission is dominated by a bright blackbody component.
In light of the hard and soft X-ray analyses of superburst and decay spectral evolution which have been reported thus far, we conclude that NuSTAR observations of a photon index of $\Gamma\approx 4$ following the 2024 superburst of MAXI J1752--457 represents a new insight into the interaction of superburst emission with the corona. While it has been shown that Type I X-ray bursts result in cooling corona (Speicher et al., 2020), superbursts offer an opportunity to study more extended interaction between burst emission and the corona. Although we did not measure the persistent emission before the burst in order to compare to our observations, given that the photon index we measured is so atypical of persistent LMXB emission, our results indicate that changes to the corona may last on the order of days after a superburst.
@@
Our analysis of the source variability did not reveal discrete features like pulsations or quasiperiodic oscillations (QPOs). Instead, we found that the source variability was dominated by red noise, which we modeled with a power law. We found a power law index of $1.8^{+0.3}_{-0.2}$, which is typical for the low-frequency noise seen in XRBs (Belloni et al., 2002), and we found that the low-frequency variability is more pronounced at higher energies. Using the Xspec tool, cpflux, we determined that the spectral power law component makes up $22\%$ of the source photon flux in the 3–6 keV range and $77\%$ of the source photon flux in the 6–30 keV range. Given a total source rms of $(13\pm 4)\%$ and $(58\pm 8)\%$ in the soft and hard bands, respectively, and if we assume that the Comptonization component is the sole source of red noise (i.e., the rms variability of the blackbody component is very small), we infer that the $10^{-3}-10^{-1}$ Hz rms of the Comptonization component was $(59\pm 18)\%$ in the 3–6 keV range and $(75\pm 10)\%$ in the 6–30 keV range. In other words, we find that the energy-dependent photon flux fraction of the spectral power law component is consistent with our measurements of the source rms as a function of energy under the assumption that the non-thermal component is the primary source of the low-frequency variability.
The significant difference in the degree of variability between the thermal and power law components raises the question of the origin of the X-ray emission during the NuSTAR observations. The thermal emission may originate from radiative cooling of the neutron star surface following the superburst, or it may be primarily due to a boundary layer heated by disk accretion (Popham & Sunyaev, 2001). Low-frequency variability is typically attributed to variable thermal disk emission for both black hole and neutron star sources (Titarchuk et al., 2007). Given that the low-frequency noise we observe can be attributed primarily to the power law spectral component, rather than to the thermal component, this would suggest that the corona is seeded by soft X-ray photons from an accretion disk such that disk variability propagates through the Comptonizing medium. Several other pieces of evidence point towards accretion powered emission. We do not observe clear evidence for cooling of the blackbody component between the two NuSTAR observations, and we find that the flux ratio between the blackbody component and the power law component did not change significantly between the two NuSTAR observations, suggesting that both the thermal and power law components were powered by accretion such that the change in accretion rate affected the overall normalization of the spectra but not the spectral shape. Furthermore, by the time that the NuSTAR observations were carried out, the source flux had become comparable to that observed by MAXI/GSC prior to the superburst. In other words, while we cannot rule out the radiative cooling scenario, we favor the interpretation that the hard X-ray emission observed by NuSTAR several days after the superburst peak was due primarily to disk accretion.
These measurements demonstrate the importance of rapid follow-up of transient sources with focusing hard X-ray observatories like NuSTAR in order to place superior constraints on the spectral evolution and physical properties of bursting sources. Further hard X-ray observations of superbursts will help to elucidate the behavior of the corona before, during, and after these long-lasting thermonuclear flashes.

### v2

@@ III.2 Timing analysis
We investigated the timing properties of MAXI J1752-457 during the second NuSTAR observation. Beginning with the clean Mode 1 and Mode 06 event files, we first corrected the photon arrival times for the motion of NuSTAR about the barycenter of the Solar System using the ftools utility, barycorr. We specified the source position as measured by Swift/XRT and reported by Liu et al. (2024). We used NuSTAR clockfile 20100101v203, produced on April 15, 2025. Next, we filtered the events using circular source regions with radii $100^{\prime\prime}$ in order to encompass as many source counts as possible given the centroid smearing in Mode 06. Unlike our spectral analysis, and in order to avoid complicated good time intervals (GTIs), we did not split the Mode 06 events into different CHU combinations.
We next used the timing analysis software packages HENDRICS and Stingray (Huppenkothen et al., 2019; Bachetti et al., 2024b, a) to complete the rest of the timing analysis. We found that the background dominates the source spectrum above 30 keV, so we filtered out events with energy below 3 keV and above 30 keV. For each focal plane module, we merged the Mode 1 and Mode 06 event lists. Using Stingray, we converted the event lists into light-curves with bins of $dt=2048^{-1}$ s. We split these light-curves into 16 segments of length $1024$ s each. We calculated the averaged power density spectrum (PDS) over these segments for FPMA and FPMB, then produced a combined PDS which we corrected for deadtime using the Fourier Amplitude Difference normalization described in Bachetti and Huppenkothen (2018). We repeated this procedure in order to produce a background PDS, using only events in background regions with radius $150^{\prime\prime}$, away from the source. The resulting Leahy-normalized PDS, binned logarithmically for clarity, are shown in the left panel of Figure 4, where the source PDS is shown in black, and the background PDS is shown in magenta. In order to investigate the energy-dependence of the source variability, we also calculated the source PDS for events with 3–6 keV and for events with 6–30 keV, shown in green and orange, respectively, in the right panel of Figure 4.
@@
IV Discussion
IV.1 Revisiting the power law index
MAXI and NinjaSat observations of MAXI J1752-457 during a superburst in November, 2024, showed that the flux and spectral evolution were consistent with an accretion-induced superburst (Aoyama et al., 2025). While these observations were sensitive to the blackbody component of the spectrum, and have demonstrated clear evolution of the temperature and radius of the blackbody, NuSTAR’s unique sensitivity above $10$ keV has allowed us to put far better constraints on the blackbody parameters following the rapid decay of the burst. We found that by the time of the second NuSTAR observation, which occured about $79$ hours after the onset of the burst, the blackbody temperature had decreased to $0.60\pm 0.1$ keV, and the inferred radius (assuming a distance of $8$ kpc) had shrunk to $6.0^{+0.4}_{-0.3}$ km.
@@ IV.1 Revisiting the power law index
NuSTAR observations also revealed that, in addition to a blackbody component, the source spectrum contained a hard X-ray excess which is well-described by a steep power law with photon index $\Gamma\approx 4$. This indicates the presence of a non-thermal plasma in which soft X-ray photons gain energy via Compton upscattering. Accreting neutron stars at similar luminosities ($10^{35-36}$$\mathrm{erg\,s^{-1}}$) tend to exhibit power law spectral components with photon indices in the range $1.5<\Gamma<3.0$ (Wijnands et al., 2015). We therefore revisit the spectra obtained during the second NuSTAR observation in order to explore the possibility of a more typical photon index.
@@
As we noted in Section III.1, when we allowed for a high-energy cutoff to be applied to the power law component, the best-fit value of $E_{\mathrm{cut}}$ was far beyond the NuSTAR energy range, indicating that a power law without a cutoff energy is preferred. However, including a high-energy cutoff allows for more flexibility in the photon index parameter as a low cutoff energy can compensate for a shallower power law. For the cutoff power law model, the $90\%$ confidence interval for the photon index is expanded, giving $\Gamma=3.7^{+0.2}_{-0.7}$. Furthermore, we calculated the $99\%$ confidence intervals for the photon index and cutoff energy by varying each parameter independently until the fit statistic had increased by 6.635. We found intervals of $2.3<\Gamma<4.0$ and $E_{\mathrm{cut}}>8$ keV. In order to explore the parameter space in even more detail, we stepped over the photon index and the cutoff energy and calculated the $W$-statistic at each step. The resulting grid is shown in Figure 5, with contours representing the 1-$\sigma$, $90\%$, and $99\%$ confidence regions. We again find that the spectra strongly prefer a steep power law with a $90\%$ confidence region that indicates a photon index greater than 2.5 and a cutoff energy above 10 keV. While photon index values of <2.5<2.5 are allowed within the 99% confidence region, they would require cutoff energies of less than 10 keV. This combination of parameters would imply a high optical depth, $\tau>7$ (Kara and García, 2025, Eq. 34). While we cannot completely rule out a more typical value for the photon index, spectral modeling of MAXI J1752-457 strongly favors an abnormally steep power law in order account for the hard X-ray excess above about 7 keV.
We note that although Aoyama et al. (2025) reported a similarly steep power law during this phase of the outburst decay, our measurement represents a distinct result for the following reasons. First, NinjaSat is sensitive to the 2-10 keV emission, whereas the observations presented here provide hard X-ray sensitivity. This is demonstrated clearly by our ability to disentangle multiple spectral components. To this point, Aoyama et al. (2025) found a steep photon index when modeling the source spectra using a power-law-only model. As we have shown, the soft X-ray is dominated by blackbody emission up to around 6 keV. Therefore the evolution of the photon index presented by Aoyama et al. (2025) should be interpreted as the softening of the blackbody component rather than a measurement of the power law component itself. Finally, where the NinjaSat spectra at this phase of the decay were limited in their constraining power — Aoyama et al. (2025) report three degrees of freedom for the fit with the steepest photon index — the observations we have presented allowed for tighter constraints and a detailed investigation of the parameter space as we have described above.
IV.2 Prior X-ray studies of superbursts
@@
Existing analyses of the hard X-ray spectra of accreting neutron stars during and after superburst are limited due to the relative infrequency of these bursts. In fact, at the time of writing, we are unable to identify any published observations of superbursts or superburst decays with NuSTAR. However, the Rossi X-ray Timing Explorer (RXTE), thanks to its All-Sky Monitor (ASM) instrument, observed 12 superbursts in the hard X-ray during its lifetime (in’t Zand, 2017). Often, analyses of the spectral evolution of superbursts with RXTE fixed the power law component to the value measured for persistent emission, typically around $\Gamma\approx 2$ (e.g. Keek et al., 2008). In some instances, significant softening of the superburst decay spectrum compared to the persistent spectrum has been observed (e.g. Kuulkers et al., 2004).
@@
Studies of superbursts in the soft X-ray band (<10<10 keV), although insensitive to the hard X-ray emission, have hinted at high photon indices during or shortly after superbursts or similar bursts. For example, Bult et al. (2021) reported on NICER measurements of the spectral evolution of IGR J17062-6143 during an intermediate duration Type I X-ray burst observed in 2020. While not categorized as a superburst, the decay of this burst lasted several days, making it comparable to the case of MAXI J1752-457. These authors reported that, when a relativistic reflection component was included in the spectral model of the decay emission, the photon index reached values in the range of $3<\Gamma<5$ throughout the two-day period following the burst before settling to more typical values around $\Gamma\approx 2.5$. Notably, however, when the spectrum was modeled without relativistic reflection, the opposite trend in the power law photon index was recovered, with $\Gamma$ increasing from around 1 to 2 throughout the burst decay.
@@
In a recently published NICER study of the 2021 superburst of 4U 1820-30, Peng et al. (2025) modeled the persistent, burst, and decay emission using a combination of a spherical blackbody and a Comptonization component represented by the Xspec model compTT. The authors noted that during the burst, the contribution of the Comptonization component could be neglected, but they reported that during the first few hours of the decay, they measured a significantly lower electron temperature than during the persistent emission, indicating a softened hard X-ray component. Jaisawal et al. (2025) presented a separate study of the same observations. These authors modeled the emission using a combination of two blackbody components and an absorbed power law. Similarly to Peng et al. (2025), these authors reported that the power law index decreased from around 2.8 back to a stable value of around 1.6 within six hours of the onset of the burst.
These NICER studies provide evidence for significant changes to non-thermal emission from accreting neutron stars during and after superbursts, but they also highlight the difficulty of disentangling spectral components with soft X-ray observations alone when the emission is dominated by a bright blackbody component. NuSTAR, on the other hand, is uniquely able to probe the hard X-ray component directly, making it possible to overcome degeneracies between the power law slope and blackbody parameters. Our result is, to the best of our knowledge, the first time that the the power law component has been constrained with hard X-ray observations in the aftermath of a superburst. The observations and results we have presented here differ from the results discussed above in other important ways. First, we have measured a steep power law index more than 5 days after the onset of the superburst in MAXI J1752-457. In contrast, Bult et al. (2021) showed that the power law index returned to more typical values within around two days after the intermediate duration Type I X-ray burst exhibited by IGR J17062-6143 in 2020, and both Peng et al. (2025) and Jaisawal et al. (2025) showed that the hard X-ray component in the spectra of 4U 1820-30 returned to a steady state within only a few hours of the 2021 superburst. Furthermore, 4U 1820-30 and IGR J17062-6143 are persistently accreting or long-term transient sources, whereas MAXI J1752-457 appears to be a transient accretor.
IV.3 Hard X-ray variability and disk-corona coupling
@@
Our analysis of the source variability did not reveal discrete features like pulsations or quasiperiodic oscillations (QPOs). Instead, we found that the source variability was dominated by red noise, which we modeled with a power law. We found a power law index of $1.8^{+0.3}_{-0.2}$, which is typical of the low-frequency noise seen in the low-luminosity state of X-ray bursters (Belloni et al., 2002), and we found that the low-frequency variability is more pronounced at higher energies. Using the Xspec tool, cpflux, we determined that the spectral power law component makes up $22\%$ of the source photon flux in the 3–6 keV range and $77\%$ of the source photon flux in the 6–30 keV range. Given a total source rms of $(13\pm 4)\%$ and $(58\pm 8)\%$ in the soft and hard bands, respectively, and if we assume that the non-thermal component is the sole source of red noise (i.e., the rms variability of the blackbody component is very small), we infer that the $10^{-3}-10^{-1}$ Hz rms of the non-thermal component was $(59\pm 18)\%$ in the 3–6 keV range and $(75\pm 10)\%$ in the 6–30 keV range. In other words, we find that the energy-dependent photon flux fraction of the spectral power law component is consistent with our measurements of the source rms as a function of energy under the assumption that the non-thermal component is the primary source of the low-frequency variability.
The significant difference in the degree of variability between the thermal and power law components raises the question of whether the source remained in the radiative cooling superburst decay phase during the NuSTAR observations, or whether the source had returned to an accretion-dominated flux state. Several pieces of evidence point towards accretion-powered emission. First, despite the fact that the two NuSTAR observations took place nearly a day apart, we do not observe evidence for cooling of the blackbody component in this interval, suggesting that the blackbody temperature had reached a persistent value. If we assume a constant blackbody radius between the first and second NuSTAR observations, we would expect to find a temperature of $kT\approx 0.53$ keV during the second observation, given the decrease in flux. We attempted to fit the spectrum observed during the second observation while fixing the blackbody radius to the value measured for the first observation, but we found that this model could not fit the data, resulting in a test statistic of $\chi^{2}/\mathrm{d.o.f.}=174/92$ (compared to $\chi^{2}/\mathrm{d.o.f.}=109/91$ when the radius was allowed to vary).
We also find that the flux ratio between the blackbody component and the power law component did not change significantly between the two NuSTAR observations, suggesting that both the thermal and power law components share a common source of energy in the form of an accretion disk. This interpretation is supported by the results of our timing analysis. Low-frequency variability is typically attributed to variable thermal disk emission for both black hole and neutron star sources (Titarchuk et al., 2007). Given that the low-frequency noise we observe can be attributed primarily to the power law spectral component, rather than to the thermal component, this would suggest that the corona is seeded by soft X-ray photons from an accretion disk such that disk variability propagates through the Comptonizing medium. Importantly, however, the flux we measured with NuSTAR remains above the pre-superburst flux. We retrieved Swift XRT observations of the source in the weeks leading up to the superburst and found that the source flux (0.3-10 keV) varied between $5\times 10^{-12}$$\mathrm{erg\,cm^{-2}\,s^{-1}}$and $5\times 10^{-11}$$\mathrm{erg\,cm^{-2}\,s^{-1}}$. On October 24, 2024, about 16 days prior to the onset of the superburst, EP measured a 0.3-10 keV flux of $(1.3\pm 0.3)\times 10^{-11}$$\mathrm{erg\,cm^{-2}\,s^{-1}}$. We also retrieved spectral data products from Swift XRT observations taken on the same day (OBSID 00016764023) and performed a simple analysis. We found that the spectrum could be described by a power law model alone ($\chi^{2}/\mathrm{d.o.f.}=192/161$), or by the combination of a spherical blackbody and a power law component ($\chi^{2}/\mathrm{d.o.f.}=183/159$). The resulting blackbody temperature and radius are $kT=0.37\pm 0.6$ keV and $R=5\pm 2$ km (assuming a distance of 8 kpc). Notably, the photon index for the combined blackbody and power law model is $\Gamma=1.1^{+1.2}_{-0.1}$. This value is more typical of accreting neutron stars than the steep photon index we observed in the NuSTAR spectra.
During the second NuSTAR observation, we measured a source flux (2-10 keV) of $6.4^{+0.3}_{-0.2}\times 10^{-11}$$\mathrm{erg\,cm^{-2}\,s^{-1}}$. The flux at this time was therefore still a factor of several times higher than the per-superburst flux. The drop in flux that we observed between the first and second NuSTAR observations is also consistent with the continued downward trend in flux observed by NinjaSat at the same time. Additionally, the pre-superburst XRT spectrum was consistent with the blackbody and power law model we used to describe the NuSTAR spectra. If we assume that accretion powers the flux in both cases, then the lower flux, lower blackbody temperature, and smaller blackbody radius we measured for the XRT spectrum is consistent with a lower rate of accretion. Importantly, however, we cannot claim that the spectra strongly prefer this model over the simple power law model.
V Conclusions
We have presented an analysis of two NuSTAR observations of MAXI J1752-457 following a superburst which occurred in November, 2024. We constrained the blackbody temperature and radius, which describe the thermal component of the X-ray spectrum, and we did not find evidence for significant cooling in the roughly 24 hours between the first and second NuSTAR observations. Timing analysis of the NuSTAR light-curve suggested that the non-thermal spectral component was seeded by photons from an accretion disk rather than from the neutron star surface. We therefore conclude that at the time of the NuSTAR observations, the source had entered a flux state powered by disk accretion rather than radiative cooling.
We also showed that the NuSTAR spectrum exhibited a steep power law tail with photon index $\Gamma\approx 4$ between 4 and 6 days after the onset of the burst, representing the first time that the power law component has been constrained with hard X-ray observations in the aftermath of a superburst. Accreting neutron stars typically exhibit power law photon indices in the range $1.5<\Gamma<3$, making our measurement of $\Gamma\approx 4$ anomalously steep. In the context of the aftermath of a superburst, this result may point towards a unique coronal state due to interaction with the burst emission. However, our ability to interpret our results in this context is limited by the lack of NuSTAR observations prior to the superburst and throughout the superburst decay. Our results demonstrate the importance of rapid follow-up and high cadence observations of superbursts with focusing hard X-ray observatories like NuSTAR in order to place superior constraints on the spectral evolution and physical properties of bursting sources. Future hard X-ray observations of superbursts will help to elucidate the behavior of the corona before, during, and after these long-lasting thermonuclear flashes.